### v2

# Distributed Perceptron under Bounded Staleness, Partial Participation, and Noisy Communication

Abstract
We study a semi-asynchronous client-server perceptron trained via iterative parameter mixing (IPM-style averaging): clients run local perceptron updates and a server forms a global model by aggregating the updates that arrive in each communication round. The setting captures three system effects in distributed deployments: (i) stale updates due to delayed model delivery and delayed application of client computations (two-sided version lag), (ii) partial participation (intermittent client availability), and (iii) imperfect communication on both downlink and uplink, modeled as effective zero-mean additive noise with bounded second moment. We introduce a server-side aggregation rule called staleness-bucket aggregation with padding that deterministically enforces a prescribed staleness profile over update ages without assuming any stochastic model for delays or participation. Under margin separability and bounded data radius, we prove a finite-horizon expected bound on the cumulative weighted number of perceptron mistakes over a given number of server rounds: the impact of delay appears only through the mean enforced staleness, whereas communication noise contributes an additional term that grows on the order of the square root of the horizon with the total noise energy. In the noiseless case, we show how a finite expected mistake budget yields an explicit finite-round stabilization bound under a mild fresh-participation condition. Synthetic experiments match the predicted square root of horizon noise scaling.
I Introduction
The perceptron is a foundational online learning algorithm with a sharp mistake bound on linearly separable data [1, 2]. Distributed variants in which clients/workers train locally and a server periodically mixes parameters were analyzed under iterative parameter mixing (IPM) [3, 4] and are structurally aligned with modern federated averaging schemes [5].
Practical distributed deployments must contend with three persistent phenomena. These effects are ubiquitous and can break classical perceptron guarantees, motivating a delay/noise-aware mistake bound handling: (i) staleness (version lag), as in bounded-staleness frameworks such as SSP [6, 7] and online learning with delayed feedback [14, 15]; (ii) partial participation, where only a subset of clients return updates each round [5, 11, 12]; and (iii) communication noise on downlink/uplink, especially in wireless settings [8, 9, 10].
We analyze a semi-asynchronous IPM-style perceptron with (i) partial participation (arbitrary arrival sets), (ii) bounded two-sided staleness (updates may be computed from older server models and applied after additional turnaround), and (iii) additive downlink/uplink noise. Our delay handling is algorithmic: staleness-bucket aggregation with padding enforces a fixed staleness profile by allocating prescribed weight to each staleness bucket when updates are available and otherwise padding with a cached historical server iterate. Our main result is a finite-horizon expected bound on the cumulative weighted number of local perceptron mistake updates: delay enters only through the enforced mean staleness, while communication noise enters through an explicit noise-energy term, yielding $O(\sqrt{A})$ growth over $A$ rounds. In the noiseless case, we further convert a finite mistake budget into a finite-round stabilization guarantee under a mild liveness condition.
II System Model and Staleness-Profile IPM Perceptron
Server rounds are $t=0,1,2,\ldots$ with server model $w_{t}\in\mathbb{R}^{D}$. An update applied at round $t$ may have been computed from a stale server model $w_{t-s}$, where $s$ is its total staleness (version lag).
We consider binary classification over $\mathbb{R}^{D}$ with labels $y\in\{-1,+1\}$. There are $m$ clients indexed by $i\in\{1,\ldots,m\}$. Client $i$ holds a finite multiset of examples $T_{i}\subseteq\mathbb{R}^{D}\times\{-1,+1\}$, and the global dataset is $T=\biguplus_{i=1}^{m}T_{i}$.
Definition 1 (Margin separability and radius).
The dataset $T$ is $\gamma$-margin separable with radius $R$ if there exist $w^{\star}\in\mathbb{R}^{D}$ with $\left\lVert w^{\star}\right\rVert=1$ and constants $\gamma>0$, $R>0$ such that
$y\,\left\langle w^{\star},\,x\right\rangle\geq\gamma,\quad\forall(x,y)\in T.$ (1)
with $\left\lVert x\right\rVert\leq R$ for all $(x,y)\in T$.
II-A Two-sided staleness and server-enforced staleness profile
The server maintains $w_{t}\in\mathbb{R}^{D}$ with initialization $w_{0}=0$ and the convention $w_{s}\equiv 0$ for all integers $s<0$. In each server round $t$, the server aggregates the set of client updates that arrive and are applied at round $t$. We denote this (possibly history-dependent) arrival set by $\mathcal{A}_{t}\subseteq\{1,\ldots,m\}$. No probabilistic model is imposed on the process $\{\mathcal{A}_{t}\}_{t\geq 0}$.
Two-sided bounded staleness
Fix bounds $\tau_{\mathrm{dl}},\tau_{\mathrm{ul}}\in\mathbb{N}_{0}$ and define the overall lag tolerance
$\tau:=\tau_{\mathrm{dl}}+\tau_{\mathrm{ul}}.$ (2)
For each arriving update $i\in\mathcal{A}_{t}$, we associate: (i) a download (downlink) staleness $s^{\mathrm{dl}}_{i,t}\in\{0,1,\ldots,\tau_{\mathrm{dl}}\}$, meaning the client starts from a server iterate that is $s^{\mathrm{dl}}_{i,t}$ rounds old at its local start time; and (ii) a turnaround delay $s^{\mathrm{ul}}_{i,t}\in\{0,1,\ldots,\tau_{\mathrm{ul}}\}$, capturing computation plus uplink latency until the update is applied at the server. Define the total staleness (version lag) at application time
$s_{i,t}:=s^{\mathrm{dl}}_{i,t}+s^{\mathrm{ul}}_{i,t}\in\{0,1,\ldots,\tau\}.$ (3)
Equivalently, if the client began local computation in round $a=t-s^{\mathrm{ul}}_{i,t}$ and read the stale model $w_{a-s^{\mathrm{dl}}_{i,t}}$, then
$w_{a-s^{\mathrm{dl}}_{i,t}}=w_{(t-s^{\mathrm{ul}}_{i,t})-s^{\mathrm{dl}}_{i,t}}=w_{t-s_{i,t}}.$ (4)
Thus the analysis depends only on the total lag $s_{i,t}$.
Client-side staleness and downlink noise
Each arriving client update $i\in\mathcal{A}_{t}$ is produced by running the perceptron on $T_{i}$ starting from a noisy stale model
$\tilde{w}^{\mathrm{dl}}_{i,t}=w_{t-s_{i,t}}+\delta_{i,t},$ (5)
where $\delta_{i,t}$ is an effective downlink noise vector. Client $i$ runs the perceptron (any order; for at least one epoch), performs $k_{i,t}$ mistake updates, and outputs a local model $w_{i,t}$.
Uplink noise
The server receives a noisy version of the returned model,
$\tilde{w}^{\mathrm{ul}}_{i,t}=w_{i,t}+\xi_{i,t},$ (6)
where $\xi_{i,t}$ is an effective uplink noise vector.
Server-enforced staleness profile via padding
Fix a target staleness profile $\bm{\alpha}=(\alpha_{0},\ldots,\alpha_{\tau})$ with $\alpha_{s}\geq 0$ and $\sum_{s=0}^{\tau}\alpha_{s}=1$. After observing the realized total stalenesses $\{s_{i,t}\}_{i\in\mathcal{A}_{t}}$, define staleness buckets
$\mathcal{B}_{s,t}:=\{i\in\mathcal{A}_{t}:\ s_{i,t}=s\},\qquad s=0,1,\ldots,\tau.$ (7)
The server assigns weights as follows:
  • If $\mathcal{B}_{s,t}\neq\emptyset$, assign nonnegative weights $\{\mu_{i,t}\}_{i\in\mathcal{B}_{s,t}}$ such that $\sum_{i\in\mathcal{B}_{s,t}}\mu_{i,t}=\alpha_{s}$ (e.g., uniform: $\mu_{i,t}=\alpha_{s}/|\mathcal{B}_{s,t}|$).
  • If $\mathcal{B}_{s,t}=\emptyset$, set a padding weight $\pi_{s,t}:=\alpha_{s}$, assigned to the cached server iterate $w_{t-s}$ (equivalently, a “virtual participant” at staleness $s$ returning $w_{t-s}$ and making 0 mistakes).
For the noisy analysis of Theorem 1, we require that $\{\mu_{i,t}\}$ and $\{\pi_{s,t}\}$ are deterministic functions of the staleness buckets $\{\mathcal{B}_{s,t}\}_{s=0}^{\tau}$ only (independent of the received model values).
The server then updates
$w_{t+1}:=\sum_{i\in\mathcal{A}_{t}}\mu_{i,t}\,\tilde{w}^{\mathrm{ul}}_{i,t}\;+\!\!\sum_{s:\,\mathcal{B}_{s,t}=\emptyset}\!\!\pi_{s,t}\,w_{t-s}.$ (8)
Protocol requirements (minimal)
The staleness-profile rule requires only that (i) the server can identify each realized total staleness $s_{i,t}$ (e.g., clients report the global model version used, or the server tags broadcasts), and (ii) the server stores the last $\tau+1$ iterates $\{w_{t-s}\}_{s=0}^{\tau}$.
Convexity
By construction all weights are nonnegative and $\sum_{i\in\mathcal{A}_{t}}\mu_{i,t}+\sum_{s:\,\mathcal{B}_{s,t}=\emptyset}\pi_{s,t}=\sum_{s=0}^{\tau}\alpha_{s}=1,$ so (8) is a convex combination.
Algorithm 1 Noisy semi-asynchronous stale-IPM perceptron with server-enforced staleness profile (padding)
1: Initialize $w_{0}\leftarrow 0$; define $w_{s}\leftarrow 0$ for all integers $s<0$.
2: Fix staleness bounds $\tau_{\mathrm{dl}},\tau_{\mathrm{ul}}\in\mathbb{N}_{0}$ and set $\tau\leftarrow\tau_{\mathrm{dl}}+\tau_{\mathrm{ul}}$.
3: Fix target staleness profile $\bm{\alpha}=(\alpha_{0},\ldots,\alpha_{\tau})$ with $\alpha_{s}\geq 0$, $\sum_{s=0}^{\tau}\alpha_{s}=1$.
4:for $t=0,1,2,\ldots$ do
5:  Server receives the set $\mathcal{A}_{t}$ of client updates that arrive and are applied at round $t$.
6:  for all $i\in\mathcal{A}_{t}$ in parallel do
7:   Server learns/infers $(s^{\mathrm{dl}}_{i,t},s^{\mathrm{ul}}_{i,t})$; set $a\leftarrow t-s^{\mathrm{ul}}_{i,t}$ and $s_{i,t}\leftarrow s^{\mathrm{dl}}_{i,t}+s^{\mathrm{ul}}_{i,t}$.
8:   Client $i$ started local computation in round $a$ from $\tilde{w}^{\mathrm{dl}}_{i,t}=w_{a-s^{\mathrm{dl}}_{i,t}}+\delta_{i,t}$ and runs local perceptron on $T_{i}$ for at least one epoch, producing $(w_{i,t},k_{i,t})$.
9:   Client returns $(w_{i,t},\,k_{i,t},\,s_{i,t})$; server receives $\tilde{w}^{\mathrm{ul}}_{i,t}=w_{i,t}+\xi_{i,t}$.
10:  end for
11:  for $s=0,1,\ldots,\tau$ do
12:   $\mathcal{B}_{s,t}\leftarrow\{i\in\mathcal{A}_{t}:\ s_{i,t}=s\}$.
13:   if $\mathcal{B}_{s,t}\neq\emptyset$ then
14:    Choose $\{\mu_{i,t}\}_{i\in\mathcal{B}_{s,t}}$ with $\mu_{i,t}\geq 0$ and $\sum_{i\in\mathcal{B}_{s,t}}\mu_{i,t}=\alpha_{s}$.
15:    (E.g.) set $\mu_{i,t}=\alpha_{s}/|\mathcal{B}_{s,t}|$ for all $i\in\mathcal{B}_{s,t}$ and set $\pi_{s,t}\leftarrow 0$.
16:   else
17:    Set padding $\pi_{s,t}\leftarrow\alpha_{s}$ on cached vector $w_{t-s}$.
18:   end if
19:  end for
20:  Update $w_{t+1}$ by staleness-profile aggregation with padding:
21:  $w_{t+1}\leftarrow\sum_{i\in\mathcal{A}_{t}}\mu_{i,t}\tilde{w}^{\mathrm{ul}}_{i,t}+\sum_{s:\,\mathcal{B}_{s,t}=\emptyset}\pi_{s,t}w_{t-s}$.
22:end for
II-B Noise model
We model communication impairments via an effective additive noise abstraction with bounded conditional second moments [10]. For each round $t$, let $H_{t}$ denote the algorithmic history up to the start of round $t$, and let $\mathcal{F}_{t}:=(H_{t},\mathcal{A}_{t},\{s_{i,t}\}_{i\in\mathcal{A}_{t}})$ denote the information available once the arrival set and stalenesses in round $t$ are revealed.
Definition 2 (Conditional zero-mean bounded-variance effective channel noise).
For every server round $t$ and every arriving client update $i\in\mathcal{A}_{t}$,
$\mathbb{E}[\delta_{i,t}\mid\mathcal{F}_{t}]$ $=0,$ $\mathbb{E}\!\left[\left\lVert\delta_{i,t}\right\rVert^{2}\mid\mathcal{F}_{t}\right]$ $\leq\sigma^{2}_{\mathrm{dl}},$ (9)
$\mathbb{E}[\xi_{i,t}\mid\mathcal{F}_{t},w_{i,t}]$ $=0,$ $\mathbb{E}\!\left[\left\lVert\xi_{i,t}\right\rVert^{2}\mid\mathcal{F}_{t},w_{i,t}\right]$ $\leq\sigma^{2}_{\mathrm{ul}}.$ (10)
Define the total per-round noise-energy bound
$V:=\sigma^{2}_{\mathrm{dl}}+\sigma^{2}_{\mathrm{ul}}.$ (11)
III Expected Weighted Mistake Bound
III-A Performance metric
Let $k_{i,t}$ be the number of local perceptron mistake updates for the client model that arrives in round $t$. Define the weighted mistakes per round
$\kappa_{t}:=\sum_{i\in\mathcal{A}_{t}}\mu_{i,t}\,k_{i,t},$ (12)
and the cumulative weighted mistakes up to horizon $A$:
$K_{A}:=\sum_{t=0}^{A-1}\kappa_{t},\qquad A\geq 1.$ (13)
(Any padding terms correspond to virtual participants with 0 mistakes.)
III-B Main theorem
Theorem 1 (Expected bound under two-sided staleness and noise).
Under margin separability (1), $\left\lVert x\right\rVert\leq R$, the bounded-staleness model (3), the noisy communication model (5)–(10), and the staleness-profile server update (8) (with $\tau=\tau_{\mathrm{dl}}+\tau_{\mathrm{ul}}$), define the mean total staleness under the profile
$\bar{s}:=\sum_{s=0}^{\tau}s\,\alpha_{s},\qquad S:=1+\bar{s}.$ (14)
Then for every horizon $A\geq 1$,
$\mathbb{E}[K_{A}]\leq\frac{SR^{2}}{\gamma^{2}}+\frac{\sqrt{SAV}}{\gamma}.$ (15)
In particular, when $V=0$ (no communication noise), (15) reduces to $\mathbb{E}[K_{A}]\leq SR^{2}/\gamma^{2}$ for all $A$.
As a corollary, (15) implies $\mathbb{E}[K_{A}]/A\to 0$ as $A\to\infty$, i.e., a vanishing expected per-round weighted mistake rate.
Remark 1 (Special cases and interpretation).
If $\tau=0$ then $S=1$ and (15) becomes $\mathbb{E}[K_{A}]\leq R^{2}/\gamma^{2}+\sqrt{AV}/\gamma$. Theorem 1 is stated in expectation; when $V=0$ the same argument gives the deterministic bound $K_{A}\leq SR^{2}/\gamma^{2}$ for any realized participation/staleness schedule. In particular, for $\tau=0$ this matches the standard IPM perceptron mistake bound [4], and for a single client it reduces to the classical perceptron bound [2]. For $\tau>0$, staleness enters only through $S=1+\bar{s}\leq\tau+1$; if $\alpha_{0}=1$ then rounds without a fresh update satisfy $w_{t+1}=w_{t}$.
IV Proof of Theorem 1
We follow the standard perceptron “progress versus norm growth” template [2] and control staleness via a staleness-augmented Lyapunov/potential argument in the spirit of asynchronous optimization analyses [13].
IV-A Single-client inequalities
Let $w^{(p)}_{i,t}$ denote client $i$’s local iterate after $p$ local mistake updates for the model that arrives in round $t$. Thus $w^{(0)}_{i,t}=\tilde{w}^{\mathrm{dl}}_{i,t}$ and for each mistake update on an example $(x^{(p)}_{i,t},y^{(p)}_{i,t})$ with $y^{(p)}_{i,t}\left\langle w^{(p)}_{i,t},\,x^{(p)}_{i,t}\right\rangle\leq 0$,
$w^{(p+1)}_{i,t}=w^{(p)}_{i,t}+y^{(p)}_{i,t}x^{(p)}_{i,t}$. (16)
The client returns $w_{i,t}:=w^{(k_{i,t})}_{i,t}$.
Lemma 1 (Single-client perceptron progress (noisy stale initialization)).
Fix a server round $t$ and $i\in\mathcal{A}_{t}$. Under (1) and $\left\lVert x\right\rVert\leq R$,
$\left\langle w^{\star},\,w_{i,t}\right\rangle$ $\geq\left\langle w^{\star},\,w_{t-s_{i,t}}\right\rangle+\left\langle w^{\star},\,\delta_{i,t}\right\rangle+\gamma k_{i,t},$ (17)
$\left\lVert w_{i,t}\right\rVert^{2}$ $\leq\left\lVert w_{t-s_{i,t}}+\delta_{i,t}\right\rVert^{2}+R^{2}k_{i,t}.$ (18)
Proof.
For any mistake update (16),
$\left\langle w^{\star},\,w^{(p+1)}_{i,t}\right\rangle$ $=\left\langle w^{\star},\,w^{(p)}_{i,t}\right\rangle+y^{(p)}_{i,t}\left\langle w^{\star},\,x^{(p)}_{i,t}\right\rangle$
$\geq\left\langle w^{\star},\,w^{(p)}_{i,t}\right\rangle+\gamma,$
by (1). Summing over $p=0,\ldots,k_{i,t}-1$ and using $w^{(0)}_{i,t}=w_{t-s_{i,t}}+\delta_{i,t}$ gives (17).
Similarly,
$\left\lVert w^{(p+1)}_{i,t}\right\rVert^{2}$ $=\left\lVert w^{(p)}_{i,t}\right\rVert^{2}+2y^{(p)}_{i,t}\left\langle w^{(p)}_{i,t},\,x^{(p)}_{i,t}\right\rangle+\left\lVert x^{(p)}_{i,t}\right\rVert^{2}$
$\leq\left\lVert w^{(p)}_{i,t}\right\rVert^{2}+R^{2},$
on a mistake, since $y^{(p)}_{i,t}\left\langle w^{(p)}_{i,t},\,x^{(p)}_{i,t}\right\rangle\leq 0$ and $\left\lVert x^{(p)}_{i,t}\right\rVert\leq R$. Summing yields (18). ∎
IV-B Global recursions in expectation
Define the scalar sequences
$a_{t}:=\left\langle w^{\star},\,w_{t}\right\rangle,\qquad b_{t}:=\left\lVert w_{t}\right\rVert^{2}.$ (19)
Lemma 2 (Stale-IPM inequalities under staleness-profile aggregation with padding).
For every server round $t\geq 0$,
$\mathbb{E}[a_{t+1}]$ $\geq\sum_{s=0}^{\tau}\alpha_{s}\,\mathbb{E}[a_{t-s}]+\gamma\,\mathbb{E}[\kappa_{t}],$ (20)
$\mathbb{E}[b_{t+1}]$ $\leq\sum_{s=0}^{\tau}\alpha_{s}\,\mathbb{E}[b_{t-s}]+R^{2}\,\mathbb{E}[\kappa_{t}]+V.$ (21)
Here we use the convention $w_{s}\equiv 0$ for $s<0$, hence $a_{s}=b_{s}=0$ for $s<0$.
Proof.
Fix a round $t$. During round $t$ a set $\mathcal{A}_{t}$ of updates arrives, with realized stalenesses $\{s_{i,t}\}_{i\in\mathcal{A}_{t}}$ and buckets $\{\mathcal{B}_{s,t}\}_{s=0}^{\tau}$.
By (8),
$w_{t+1}=\sum_{i\in\mathcal{A}_{t}}\mu_{i,t}(w_{i,t}+\xi_{i,t})\;+\!\!\sum_{s:\,\mathcal{B}_{s,t}=\emptyset}\!\!\pi_{s,t}\,w_{t-s}.$ (22)
Progress recursion
Taking $\left\langle w^{\star},\,\cdot\right\rangle$ in (22) and applying Lemma 1 gives
$a_{t+1}$ $\geq\sum_{i\in\mathcal{A}_{t}}\mu_{i,t}\,a_{t-s_{i,t}}+\gamma\,\kappa_{t}$
$\quad+\sum_{i\in\mathcal{A}_{t}}\mu_{i,t}\left\langle w^{\star},\,\delta_{i,t}\right\rangle+\sum_{i\in\mathcal{A}_{t}}\mu_{i,t}\left\langle w^{\star},\,\xi_{i,t}\right\rangle$
$\quad+\sum_{s:\,\mathcal{B}_{s,t}=\emptyset}\pi_{s,t}a_{t-s}.$
Condition on $\mathcal{F}_{t}$. Under this conditioning, the weights are fixed (they depend only on buckets) and the noise terms have zero conditional mean given $\mathcal{F}_{t}$ (for $\xi_{i,t}$ this follows from (10) by iterated expectation over $w_{i,t}$), hence their conditional expectations vanish. Finally, by construction of the profile, for any scalar sequence $\{Z_{t-s}\}_{s=0}^{\tau}$,
$\sum_{i\in\mathcal{A}_{t}}\mu_{i,t}Z_{t-s_{i,t}}+\sum_{s:\,\mathcal{B}_{s,t}=\emptyset}\pi_{s,t}Z_{t-s}=\sum_{s=0}^{\tau}\alpha_{s}Z_{t-s}.$ (23)
Applying (23) with $Z_{t-s}=a_{t-s}$ and then using the law of total expectation yields (20).
Norm recursion
Since (22) is a convex combination, Jensen’s inequality gives
$b_{t+1}=\left\lVert w_{t+1}\right\rVert^{2}\leq\sum_{i\in\mathcal{A}_{t}}\mu_{i,t}\left\lVert w_{i,t}+\xi_{i,t}\right\rVert^{2}+\sum_{s:\,\mathcal{B}_{s,t}=\emptyset}\pi_{s,t}b_{t-s}.$
Condition on $\mathcal{F}_{t}$. Using (10),
$\mathbb{E}[\left\lVert w_{i,t}+\xi_{i,t}\right\rVert^{2}\mid\mathcal{F}_{t}]\leq\mathbb{E}[\left\lVert w_{i,t}\right\rVert^{2}\mid\mathcal{F}_{t}]+\sigma^{2}_{\mathrm{ul}}.$
By Lemma 1, $\left\lVert w_{i,t}\right\rVert^{2}\leq\left\lVert w_{t-s_{i,t}}+\delta_{i,t}\right\rVert^{2}+R^{2}k_{i,t}$, and (9) implies $\mathbb{E}[\left\lVert w_{t-s_{i,t}}+\delta_{i,t}\right\rVert^{2}\mid\mathcal{F}_{t}]\leq b_{t-s_{i,t}}+\sigma^{2}_{\mathrm{dl}}$. Combining these bounds and substituting into Jensen yields
$\begin{split}\mathbb{E}[b_{t+1}\mid\mathcal{F}_{t}]&\leq\sum_{i\in\mathcal{A}_{t}}\mu_{i,t}b_{t-s_{i,t}}+\sum_{s:\,\mathcal{B}_{s,t}=\emptyset}\pi_{s,t}b_{t-s}\\ &\quad+R^{2}\,\mathbb{E}[\kappa_{t}\mid\mathcal{F}_{t}]+V,\end{split}$
where $V=\sigma^{2}_{\mathrm{dl}}+\sigma^{2}_{\mathrm{ul}}$. Applying (23) with $Z_{t-s}=b_{t-s}$ and then using the law of total expectation yields (21).
IV-C Lyapunov potential and conclusion
Define tail sums $c_{0}:=1$ and for $j\in\{1,\ldots,\tau\}$,
$c_{j}:=\sum_{s=j}^{\tau}\alpha_{s}.$ (24)
Define the potentials for $t\geq 0$:
$\Phi_{t}:=\sum_{j=0}^{\tau}c_{j}\,\mathbb{E}[a_{t-j}],\qquad\Psi_{t}:=\sum_{j=0}^{\tau}c_{j}\,\mathbb{E}[b_{t-j}].$ (25)
Using $c_{j}-c_{j+1}=\alpha_{j}$ (with $c_{\tau+1}:=0$), one can rewrite $\Phi_{t+1}=\Phi_{t}+\mathbb{E}[a_{t+1}]-\sum_{s=0}^{\tau}\alpha_{s}\mathbb{E}[a_{t-s}]$ (and similarly for $\Psi_{t}$); substituting Lemma 2 yields the one-step bounds
$\Phi_{t+1}\geq\Phi_{t}+\gamma\,\mathbb{E}[\kappa_{t}],\qquad\Psi_{t+1}\leq\Psi_{t}+R^{2}\,\mathbb{E}[\kappa_{t}]+V.$ (26)
Telescoping from $t=0$ to A1A-1 gives
$\Phi_{A}\geq\gamma\,\mathbb{E}[K_{A}],\qquad\Psi_{A}\leq R^{2}\,\mathbb{E}[K_{A}]+AV.$ (27)
Let $Z:=\sum_{j=0}^{\tau}c_{j}w_{A-j}$. By linearity and $\left\lVert w^{\star}\right\rVert=1$, $\Phi_{A}=\mathbb{E}[\left\langle w^{\star},\,Z\right\rangle]\leq\mathbb{E}[\left\lVert Z\right\rVert]$. Moreover, using Jensen and Cauchy–Schwarz (with $c_{j}\geq 0$),
$\mathbb{E}[\left\lVert Z\right\rVert]^{2}$ $\leq\mathbb{E}[\left\lVert Z\right\rVert^{2}]$
$\leq\left(\sum_{j=0}^{\tau}c_{j}\right)\,\mathbb{E}\!\left[\sum_{j=0}^{\tau}c_{j}\,\left\lVert w_{A-j}\right\rVert^{2}\right]$
$=\left(\sum_{j=0}^{\tau}c_{j}\right)\Psi_{A}.$
A direct calculation yields
$\sum_{j=0}^{\tau}c_{j}=1+\sum_{s=0}^{\tau}s\,\alpha_{s}=S.$ (28)
Thus $\Phi_{A}^{2}\leq S\Psi_{A}$. Combining with (27) gives $\gamma^{2}\mathbb{E}[K_{A}]^{2}\leq S(R^{2}\mathbb{E}[K_{A}]+AV)$; solving for $\mathbb{E}[K_{A}]$ yields (15).
V Finite-Round Stabilization in the Noiseless Case
Figure 1: Synthetic illustration of Theorem 1 for Algorithm 1. (a) Noiseless links: different enforced staleness profiles $\bm{\alpha}$ (legend shows $\bar{s}$) lead to different cumulative weighted mistake levels. (b) Fixed profile: increasing noise energy $V$ increases $\mathbb{E}[K_{A}]$ approximately as $C+c\sqrt{A}$.
Theorem 1 controls cumulative (weighted) mistakes. In general, a finite mistake bound does not necessarily imply a finite-round bound in semi-asynchronous systems: counterexamples can exist. We show that a finite-round stabilization bound does follow in the noiseless case ($V=0$), under a mild liveness assumption and a simple within-bucket weighting property.
Throughout this section we specialize to noiseless links:
$\delta_{i,t}\equiv 0,\qquad\xi_{i,t}\equiv 0,$ (29)
and define $K_{\infty}:=\sum_{t=0}^{\infty}\kappa_{t}$.
Theorem 2 (Finite-round stabilization in the noiseless case ($V=0$)).
Assume (29) and $\alpha_{0}>0$, and that every participating client executes at least one complete pass over its local dataset $T_{i}$ in Algorithm 1. Assume further:
  • (L1)
    (Uniform fresh participation) There exists $p_{\min}>0$ such that for every client $i$ and every round $t$,
    $\mathbb{P}\big(i\in\mathcal{A}_{t},\ s_{i,t}=0\ \big|\ H_{t}\big)\ \geq\ p_{\min}$
The server uses any within-bucket rule (in the noiseless setting this may depend on the reported mistake counts $\{k_{i,t}\}$) such that whenever some fresh client makes a mistake in round $t$,
$\sum_{i\in\mathcal{B}_{0,t}}\mu_{i,t}k_{i,t}\ \geq\ \alpha_{0}.$
Define
$\begin{split}T_{\mathrm{hit}}&:=\inf\{t\geq 0:\ y\left\langle w_{t},\,x\right\rangle>0,\ \forall(x,y)\in T\},\\ T_{\mathrm{stab}}&:=\inf\{t\geq 0:\ y\left\langle w_{r},\,x\right\rangle>0,\ \forall(x,y)\in T,\ \forall r\geq t\}.\end{split}$ (30)
Then
$\mathbb{E}[T_{\mathrm{hit}}]$ $\leq\frac{SR^{2}}{\alpha_{0}p_{\min}\gamma^{2}},$ (31)
$\mathbb{E}[T_{\mathrm{stab}}]$ $\leq(\tau+1)\frac{SR^{2}}{\alpha_{0}p_{\min}\gamma^{2}}.$
Remark 2 (Noisy vs. noiseless within-bucket weighting).
In Theorem 1 (noisy links), the bucket-only determinism of $\{\mu_{i,t}\}$ is used in Lemma 2 so that, conditioning on $\mathcal{F}_{t}$, the weights are fixed and the zero-mean noise terms vanish in expectation. Under (29) (i.e., $V=0$), these noise terms are identically zero, so within-bucket weights may be chosen in any (possibly mistake-aware) way to enforce the fresh-mistake weight condition; e.g., put the fresh mass $\alpha_{0}$ uniformly on $\{i\in\mathcal{B}_{0,t}:k_{i,t}>0\}$ (and zero on the rest).
Proof.
Work in the noiseless regime (29). For each round $t$, define the event
$\mathcal{E}_{t}:=\{\,y\left\langle w_{t},\,x\right\rangle\leq 0\text{ for some }(x,y)\in T\,\},$
i.e., $\mathcal{E}_{t}$ is the event that $w_{t}$ is not globally correct.
Step 1: A one-round lower bound on $\mathbb{E}[\kappa_{t}\mid H_{t}]$.
Fix $t\geq 0$ and condition on $H_{t}$.
If $\mathcal{E}_{t}$ does not occur then $\mathbf{1}\{\mathcal{E}_{t}\}=0$ and the inequality below is trivial. Otherwise, since $T=\biguplus_{i=1}^{m}T_{i}$, there exists a client index $i_{t}\in\{1,\ldots,m\}$ and some $(x,y)\in T_{i_{t}}$ such that $y\left\langle w_{t},\,x\right\rangle\leq 0$.
Consider the fresh-participation event
$G_{t}:=\{\,i_{t}\in\mathcal{A}_{t},\ s_{i_{t},t}=0\,\}.$
By (L1), $\mathbb{P}(G_{t}\mid H_{t})\geq p_{\min}$. On $G_{t}$, client $i_{t}$ starts from $w_{t}$ (noiseless downlink) and executes at least one epoch over $T_{i_{t}}$. Because $w_{t}$ misclassifies at least one example in $T_{i_{t}}$, the first epoch must contain at least one perceptron mistake update, hence $k_{i_{t},t}\geq 1$. Therefore, on $G_{t}$ there exists a fresh client that makes a mistake in round $t$, and by the fresh-mistake weight condition we have
$\kappa_{t}=\sum_{i\in\mathcal{A}_{t}}\mu_{i,t}k_{i,t}\geq\sum_{i\in\mathcal{B}_{0,t}}\mu_{i,t}k_{i,t}\geq\alpha_{0}.$
Consequently,
$\mathbb{E}[\kappa_{t}\mid H_{t}]\geq\alpha_{0}p_{\min}\,\mathbf{1}\{\mathcal{E}_{t}\}.$
Step 2: Bounding the expected number of incorrect rounds.
Taking expectations and summing over $t\geq 0$ gives
$\mathbb{E}[K_{\infty}]=\sum_{t=0}^{\infty}\mathbb{E}[\kappa_{t}]=\sum_{t=0}^{\infty}\mathbb{E}\big[\mathbb{E}[\kappa_{t}\mid H_{t}]\big]\geq\alpha_{0}p_{\min}\sum_{t=0}^{\infty}\mathbb{P}(\mathcal{E}_{t}).$
By Theorem 1 with $V=0$, $\mathbb{E}[K_{A}]\leq SR^{2}/\gamma^{2}$ for all $A$, hence $\mathbb{E}[K_{\infty}]\leq SR^{2}/\gamma^{2}$. Therefore
$\sum_{t=0}^{\infty}\mathbb{P}(\mathcal{E}_{t})\leq\frac{SR^{2}}{\alpha_{0}p_{\min}\gamma^{2}}.$
Bound for $T_{\mathrm{hit}}$.
By definition of $T_{\mathrm{hit}}$ in (30), for every sample path and every $t<T_{\mathrm{hit}}$ the iterate $w_{t}$ is not globally correct, i.e., $\mathcal{E}_{t}$ occurs. Hence
$T_{\mathrm{hit}}=\sum_{t=0}^{\infty}\mathbf{1}\{t<T_{\mathrm{hit}}\}\leq\sum_{t=0}^{\infty}\mathbf{1}\{\mathcal{E}_{t}\}.$
Taking expectations yields
$\mathbb{E}[T_{\mathrm{hit}}]\leq\sum_{t=0}^{\infty}\mathbb{P}(\mathcal{E}_{t})\leq\frac{SR^{2}}{\alpha_{0}p_{\min}\gamma^{2}}.$
A window condition implying stabilization.
Suppose that for some round $t$ we have $y\left\langle w_{t-j},\,x\right\rangle>0$ for all $(x,y)\in T$ and all $j=0,1,\ldots,\tau$. Consider any client update applied at round $t$: its total staleness satisfies $s_{i,t}\leq\tau$, so the client starts from $w_{t-s_{i,t}}$, which is globally correct and in particular correct on $T_{i}$. Under Algorithm 1, this implies the client makes no local mistake updates and returns $w_{i,t}=w_{t-s_{i,t}}$. Padding terms in (8) are also among $\{w_{t-j}\}_{j=0}^{\tau}$. Therefore $w_{t+1}$ is a convex combination of globally correct vectors. Since the set $\{w:\ y\left\langle w,\,x\right\rangle>0,\ \forall(x,y)\in T\}$ is an intersection of (open) halfspaces, it is convex, hence $w_{t+1}$ is also globally correct. Repeating the same argument inductively shows $w_{r}$ is globally correct for all $r\geq t$, i.e., $T_{\mathrm{stab}}\leq t$.
Equivalently, if $t<T_{\mathrm{stab}}$ then the window $\{w_{t-j}\}_{j=0}^{\tau}$ cannot be entirely correct, so at least one of $w_{t},w_{t-1},\ldots,w_{t-\tau}$ is incorrect. Thus, with the convention $\mathbf{1}\{\mathcal{E}_{s}\}=0$ for $s<0$,
$\mathbf{1}\{t<T_{\mathrm{stab}}\}\leq\sum_{j=0}^{\tau}\mathbf{1}\{\mathcal{E}_{t-j}\}.$
Summing over $t\geq 0$ gives
$T_{\mathrm{stab}}$ $=\sum_{t=0}^{\infty}\mathbf{1}\{t<T_{\mathrm{stab}}\}$
$\leq\sum_{t=0}^{\infty}\sum_{j=0}^{\tau}\mathbf{1}\{\mathcal{E}_{t-j}\}$
$=(\tau+1)\sum_{t=0}^{\infty}\mathbf{1}\{\mathcal{E}_{t}\}.$
Taking expectations and using the bound on $\sum_{t\geq 0}\mathbb{P}(\mathcal{E}_{t})$ yields
$\mathbb{E}[T_{\mathrm{stab}}]\leq(\tau+1)\sum_{t=0}^{\infty}\mathbb{P}(E_{t})\leq(\tau+1)\frac{SR^{2}}{\alpha_{0}p_{\min}\gamma^{2}}.$
This establishes (31). ∎
VI Experiments
We provide a synthetic experiment to illustrate the qualitative behavior predicted by Theorem 1. We generate linearly separable data satisfying Definition 1, partition it across clients, and run Algorithm 1 for $A$ server rounds under partial participation and bounded total staleness $\tau=\tau_{\mathrm{dl}}+\tau_{\mathrm{ul}}$. For illustration we use a simple randomized participation/delay schedule (the theory does not require any such model) and evaluate several enforced profiles $\bm{\alpha}$; Fig. 1 reports Monte Carlo estimates of $\mathbb{E}[K_{A}]$. In the noiseless case ($V=0$), panel (a) shows that emphasizing fresher updates (smaller enforced mean staleness $\bar{s}$) yields a smaller mistake budget, consistent with the $S=1+\bar{s}$ dependence in (15). Panel (b) fixes $\bm{\alpha}$ and varies the total noise energy $V$ (implemented via zero-mean additive perturbations with bounded second moment), illustrating the predicted $O(\sqrt{A})$ growth in (15).
VII Discussion
Algorithm 1 implements a deterministic age-mixing rule: in every round the server forms a convex combination whose total-staleness distribution matches a chosen profile $\bm{\alpha}$, padding missing buckets with cached iterates. This removes the need to model delays or participation and explains Theorem 1: asynchrony enters only through the enforced mean staleness $\bar{s}$ (via $S=1+\bar{s}$), while link noise contributes through $V$ in (15). Since $S$ appears linearly in $SR^{2}/\gamma^{2}$ and inside the square root in $\sqrt{SAV}/\gamma$, reducing $\bar{s}$ improves both the noiseless plateau and the noisy $\sqrt{A}$ slope.
A practical profile-design heuristic is to estimate bucket-occupancy frequencies online and choose $\bm{\alpha}$ supported on stalenesses that are reliably present; otherwise that mass is frequently assigned to cached iterates, which is safe but reduces the fraction of new information mixed each round. Among feasible profiles, (15) suggests minimizing $\bar{s}$ to shrink both terms. Finally, (15) also reveals a noise-limited regime: once $\sqrt{SAV}/\gamma$ dominates $SR^{2}/\gamma^{2}$, additional rounds yield diminishing returns unless the effective $V$ is reduced.
VIII Conclusion
We analyzed a semi-asynchronous distributed perceptron with bounded downlink staleness, bounded uplink/computation turnaround delay, partial participation, and additive uplink/downlink communication noise. The finite-horizon guarantee in Theorem 1 makes the dependence on the server-enforced mean total staleness and on noise energy explicit, and recovers standard perceptron [2] and IPM [4] results as special cases. In addition, in the noiseless regime, Theorem 2 yields a finite-round stabilization guarantee under a mild fresh-participation condition.